OpenStudy (anonymous):

Solve this linear system in 2D [x, y]= [-12, -7] + s [8, -5] [x, y]= [2, -1] + t [3, -2]

OpenStudy (anonymous):

is it not -12+8s = 2 + 3t and -7 -5s =-1-2t and solve for s, t and then back to x,y?

OpenStudy (anonymous):

subs s and t into the original ? @Hoa

OpenStudy (anonymous):

I think so,

OpenStudy (anonymous):

from the derive system, solve for s and t and then put them back to original one to get x, y.

OpenStudy (anonymous):

okay thanks !

OpenStudy (anonymous):

np

OpenStudy (anonymous):

What do i do after substituting it into (1) and (2) ? @Hoa

OpenStudy (anonymous):

sorry, I thought you get it so I left for others. after getting s and t, you replace to one of the original [x,y] to get values of x and y . conclude that [x,y] = [ number you get, number you get] . that's it

OpenStudy (anonymous):

I got s =46, now replace [x,y] = [-12,-7] + 46[8, -5] = [356, -237] I think so, if it's not the answer, I am sorry, I cannot help